Suppose that Total revenue of the firm is
TR = 5Q + 3Q2 and TC is TC = Q + 4Q2.
Obtain the quantity that maximizes profit and hence maximum profit.
1
Expert's answer
2019-11-08T06:24:50-0500

Maximum profit occurs when:


"MR = MC"

and


"MR = 5 + 6Q""MC = 1+ 8Q"

For maximum profit:

"5+6Q =1+8Q ;"

"5-1 =8Q-6Q"

"2Q =4"

"\therefore Q=2"


Maximum profit = TR- TC



"5(10)+3(2^2 )- (2+4(2^2))="


"62-18=44""Max Profit = 44"